OpenStudy (keyonsay):

Simplify the following expression. 14/15 - 1/4=

OpenStudy (anonymous):

Find the common denominator of 15 and 4: 60 \[\frac{ 14 }{ 15 }=\frac{ 56 }{ 60 }, \frac{ 1 }{ 4 }=\frac{ 15 }{ 60 }\] \[\frac{ 56 }{ 60 }-\frac{ 15 }{ 60 }=\frac{ 41 }{ 60 }\] So, \[\frac{ 14 }{ 15 }-\frac{ 1 }{ 4 }=\frac{ 41 }{ 60 }\]

OpenStudy (keyonsay):

you sure are helping me alot :DDD thank you

OpenStudy (anonymous):

:D No problem!